A stone is released at the top of a 1.2 m tall barrel filled with liquid. The stone reaches a speed of 1.7 m/s just before it hits the bottom of the barrel 0.90 s later. Assuming constant acceleration, find the stone’s average speed and average velocity


1
Expert's answer
2021-03-24T19:38:35-0400

Average velocity is the average of initial and final velocity:


"v=\frac{v_i+v_f}{2}=\frac{0+1.7}{2}=0.85\text{ m/s}."

Average speed is total distance divided by total time:


"s=\frac{1.2}{0.9}=1.33 \text{ m/s}."